FEL Upgrade Project Weekly Brief
December 16-20, 2002


Message to our Readers:
We thank all of our stakeholders and collaborators who have helped us design, build, install and plan for the use of the FEL Upgrade over the past year and who will continue to join us as we bring the device into operation in 2003. The past year has been an extremely exciting, productive and exhausting year for all of us involved in the project. We wish all of you an enjoyable and safe holiday break with your friends.

Highlights for the Week:
The team worked hard to bring many installation, check-out and commissioning activities to a point where our operations can be shut down for our annual holiday break. (Jefferson Lab will be officially closed from Dec. 21 through January 1. Our next report will come to you on Friday January 10th.) Highlights of the week’s activities include: (1) first check-out of the controls of the optical cavities; (2) operation of the complete RF system for the injector to 100kW (the full required power); (3) all cavity assemblies for the 3rd cryomodule were completed ahead of schedule and we are ready for cold RF tests next month (kudos to the SRF team); (4) the first GX (reverse bend) dipole magnet on the test stand has met its demanding (1/10,000) field flatness specification (kudos to the Magnetic Measurement Team); (5) the first arc 180 degree dipole magnet (DY) is ready for installation after tightening of the top plate assembly over the newly installed modified vacuum chamber.

Special Note Concerning Postponement of January 23-24 FEL Users/LPC Meeting:

Our annual FEL Users Meeting and LPC Workshop, which was scheduled for January 23-24, 2003 has been postponed to March 12-14, 2003 in order to combine this event with a planning workshop on potential scientific applications of the FEL Upgrades. Further information and the agenda for this combined meeting will be distributed in this note and on our web site in January.

Due to the very tight FY03 budget situation, we are now preparing weekly financial status reports and sending them to DOE and ONR. The first report was sent last Friday and today’s report will be followed by a report on Jan.10th , which will summarize the first full week of January.

Project Monthly reports for October and November were completed. We fell behind in these routine reports due to the extra work load of preparing special cost and schedule reports for DOE and ONR over the last two months.

SURA and DOE are planning a review of the potential scientific applications of the 10 kW IR Upgrade, 1 kW UV Upgrade and subsequent potential upgrades (100 kW IR) for the spring of 2003. A workshop is being planned for March 12-14th, coincident with the annual FEL Users/LPC workshop on March 12, to plan for the review which may be scheduled for June 2003.

WBS 4 (Injector):
We are in still in the process of baking the gun and downstream systems. We had a 1-day stop when the system reached 200C to examine the residual gas behavior. There some concern that we had opened up a leak as the system moved from the previous bakeout temperature of 150C to 200C. However, the high levels at 28amu we were seeing were due to a mixture of N2 outgassing from the electrodes, CO degassing from our getter pumps and the residual hydrocarbon contamination. The hydrocarbon levels are decreasing but still too high for our liking. We are pumping into a turbo pump and will shift to internal ion pumps when the levels drop sufficiently. We will keep the system baking until just before Dec.25th when we will give our intrepid Mr. Clean (Tim Siggins) a well-deserved holiday break.

Gun HVPS - The customer interface box was operated successfully with the HVPS under EPICS control. Several readouts still need calibration factors, but they are relatively minor changes. The HVPS correctly provides the requested voltage subject to the current limit setting.

WBS 5 (SRF):
All cavity assemblies (eight plus one spare) were completed for the 3rd cryomodule ahead of schedule. Half the batch has been tuned and aligned. Therefore, we are ready for cold RF tests in the Vertical Test Area (VTA) next month. We decided to weld all helium vessels in place before VTA testing to save one cycle in the testing schedule, following the scheme used for SNS cavities.

WBS 6 (RF):
Quarter RF - Quarter klystrons were operated to 100 kW into waveguide shorts. This tested the klystrons, HVPS, circulators, RF loads, and software for both systems. The final correction values will be checked today. SRF commissioning can start in January.

Buncher - The Buncher is presently disassembled for the Gun bake.

Zone 2 - Ready for operations.

Zone 4 - SRF commissioning can be completed in January.

WBS 8 (Instrumentation):
Software for the 3F and 4F trim magnets software has been completed, 1st, 2nd Arc, Wiggler, Injection and extraction shunts modules software has also been completed. Work is still in progress for the respective Ops screens. Thursday was spent with the folks at BNL reviewing hardware and software for the Danfisik super trim power supplies. On Friday we were able to begin the testing through the serial port (laptop) on the PSC VME board. The re-alignment for 0F06 sub region has been completed and is awaiting re-assembly of the trim magnets in order to complete high level checkout of the 0F region.

Fabrication of 2 BV interface boxes is in progress for the ultra viewers. High level checkout of the beam viewers in the 0F to 3F09 is in progress. This checkout includes repeatability, alignment, brightness, EPICS interface, Ops screen. Installation of IXM0F06 (Terra hertz mirror) is complete (done, done). Five (5) stepper motor front panel modules were received and are being populated in house. Wiring of the stepper motor chassis will be completed as the front panel modules are populated.

The high voltage power supply controls are all up and running. The digital PCB artwork was sent out for the production boards. These are due Dec. 27. The current board was hand wired. The update to the analog control card are being done now and will be sent out when we return from holiday.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
• The DY magnet was assembled, but for placing the top by the folks at Magnet test. Don Bullard, who was scheduled to put it together, has not been available because of extended bakeout of the vacuum system on the injector gun. The DY will be fully assembled on January 2 and shipped over to the enclosure.
Diversion of the beam after the Wiggler.
• The vacuum group is working on the vacuum pipes and the stand in the extraction area beyond the wiggler.
• The GW magnet is placed on its stand and hook up has started.
Optical Chicane Dipoles (GW)
• Magnet Test will do testing on these UV Line magnets after the IR Magnets are fully tested.
Injector Dipoles (DU/DV)
• No action
Arc 180 Degree Dipoles (GY)
• Wang NMR completed the impregnation of the second coil and cured it. I will visit them on Monday in order to make sure assembly is correct. It is due to ship by Jan. 2.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
• The first GX obtained David Douglas’ ± 1 part in 10,000 criteria for core field at all three excitations and over a wider good field region than required. It also matches the GQ within 1 part in 1000 in field at a known current. The magnet is out of specification in two criteria. Fortunately we have knobs and handle that will bring the magnet into specification.

The first miss is that the integral is high by 1.3%. Fortunately, this magnet is a pretty severe wedge and has a wider good field region. We can be move it transverse to the beam by the 0.6 inches necessary to bring the field integral within specification.

The next miss is that the angle of the magnet is off by 0.2 %. We believe this is a result of the 210 MeV pole face shims that are installed during these runs. We will remove them and alter the angle of the field clamps to the pole end faces to bring this property within specification.

• Wang is ready to ship the remaining two GX magnets with the first GY and is working on the GQ coils for a mid January shipment of those remaining magnets.
Trim Quad (QT)
• The trim coil to create the GC vertical corrector function within the trim quad is out for bid.
Sextupole (SF)
• No progress
Octupole (OT)
• The procurement process continues.
Beam Line and Vacuum
• The vacuum group continued leak checking and qualifying the remaining chambers for the Arcs. All are on site.
• Optical Chicane chamber fabrication in the JLab Shop continued. Three of the required four are complete.

WBS 11 (Optics):
Activities this week emphasized putting the high reflector cavity hardware through it's paces while under vacuum. We moved these activities to the front of the queue in order to (1) get the Vespel leadscrew nut to lubricate the leadscrew, which must be done initially under vacuum, and (2) test the cooling loop conveyor chain. This was our first opportunity to connect our installed hardware to the hardware (cabling and motion controls) installed by the I&C group, and overall, it went pretty well. In the process we gained experience in the techniques we need to master for working on the cavity hardware in the FEL vault, rather than upstairs. At this time (morning of 12/20), the HR pressure is still too high to operate the leadscrew. The primary component of the pressure is adsorbed water, and we are assessing which technique to use to remove it.

Fabrication of the diagnostic viewer/laser safety shutter and the insertable mirrors continues. The insertable mirror hardware that will be used for injector e-beam studies was installed and is being aligned.

Other activities:
The laser oscillator for the JTO-funded drive laser project was awarded to Polytec PI/Time Bandwidth Products. We learned that the laser system (IR tunable, ps pulsewidth) acquired by Prof. Gunter Luepke will arrive at the beginning of the new year. This laser system will be installed in User Lab 6, and will be a versatile probe synchronized to the FEL.